# Shallow IQP circuits and graph generation

### v2

@@ Abstract
We introduce shallow instantaneous quantum polynomial-time (IQP) circuits as generative graph models, using an edge-qubit encoding to map graphs onto quantum states. Focusing on bipartite and Erdős–Rényi distributions, we study their expressivity and robustness through simulations and large-scale experiments. Noiseless simulations of $28$ qubits (88-node graphs) reveal that shallow IQP models can learn key structural features, such as the edge density and bipartite partitioning. On IBM’s Aachen QPU, we scale experiments from $28$ to $153$ qubits (881818 nodes) in order to characterize performance on real quantum hardware. Local statistics—such as the degree distributions—remain accurate across scales with total variation distances ranging from $0.04$ to $0.20$, while global properties like strict bipartiteness degrade at the largest system sizes ($91$ and $153$ qubits). Notably, spectral bipartivity—a relaxation of strict bipartiteness—remains comparatively robust at higher qubit counts. These results establish practical baselines for the performance of shallow IQP circuits on current quantum hardware and demonstrate that, even without error mitigation, such circuits can learn and reproduce meaningful structural patterns in graph data, guiding future developments in quantum generative modeling for the NISQ era and beyond.
@@ I Introduction
Graphs represent relationships between entities and provide a unified framework to describe complex systems via their components and interactions [newmanNetworks2018]. Their versatility has driven the development of graph generation methods for applications ranging from drug discovery to scheduling [drobyshevskiyRandomGraphModeling2020, bonifatiGraphGeneratorsState2021, bonginiMolecularGenerativeGraph2021, yangMoleculeGenerationDrug2024, cordeiroRandomGraphGeneration2010].
Machine learning (ML) methods have emerged as powerful tools for graph generation, mirroring their success in other generative modeling tasks [guoSystematicSurveyDeep2022]. These approaches typically fall into two groups [guoSystematicSurveyDeep2022]: sequential methods, which build graphs autoregressively by adding nodes, edges, or motifs [groverGraphiteIterativeGenerative2019, tranDeepNCDeepGenerative2020, bacciuEdgebasedSequentialGraph2020], and one-shot methods, which generate entire graphs in a single pass by sampling from a learned distribution [simonovskyGraphVAEGenerationSmall2018, flam-shepherdGraphDeconvolutionalGeneration2020, niuPermutationInvariantGraph2020]. Sequential methods provide finer control over generation but incur higher sampling costs; one-shot methods trade control for speed.
@@
Nevertheless, learning a graph distribution remains challenging because node interactions span both local and global dependencies. As the number of nodes increases, the combinatorial complexity of pairwise and higher-order relationships grows rapidly: adding a single node to an $M$-node graph introduces $M$ new potential edges, so that the number of possible configurations increases from $2^{\binom{M}{2}}$ to $2^{\binom{M}{2}}\cdot 2^{M}$. This exponential scaling implies that even modest increases in graph size can dramatically expand the hypothesis space. Although practical problems typically restrict attention to structured graph families, models must still possess sufficient expressivity to navigate this high-dimensional space.
Quantum generative modeling emerges as an alternative framework to overcome the limitations of its classical counterpart by leveraging quantum phenomena [zoufalGenerativeQuantumMachine2021]. By encoding information in quantum states, quantum generative models (QGMs) operate in a Hilbert space whose dimension grows exponentially with the number of qubits [nielsenQuantumComputationQuantum2010], allowing them to represent probability distributions that classical systems cannot feasibly store or manipulate. Consequently, QGMs can efficiently learn and reproduce distributions that are intractable for classical models [swekeQuantumClassicalLearnability2021, huangGenerativeQuantumAdvantage2025]. Whether such beyond-classical expressivity yields practical benefits for structured domains such as graph generation remains an open question.
A prominent example of a QGM is the quantum circuit Born machine (QCBM), which encodes probability distributions via the measurement statistics of a parameterized quantum circuit [coyleBornSupremacyQuantum2020]. QCBMs can be implemented using an instantaneous quantum polytime (IQP) ansatz [shepherdInstantaneousQuantumComputation2009], an architecture that has emerged as a promising candidate for quantum generative modeling [recio-armengolTrainClassicalDeploy2025, kurkinUniversalityKerneladaptiveTraining2025, zouConQuERModularArchitectures2025]. These circuits are conjectured to be classically intractable [marshallImprovedSeparationQuantum2024], yet certain statistical properties of their output distributions can be efficiently estimated on classical hardware. This feature enables a hybrid learning framework where training relies on classical estimated statistics, while sampling requires quantum hardware [recio-armengolTrainClassicalDeploy2025].
In this work, we investigate shallow IQP QCBMs for generating random bipartite and Erdős–Rényi graph distributions. These families offer an ideal testbed: their well-defined combinatorial structure and varied features probe different facets of a model’s expressive power. We propose a single-shot graph-generation workflow that trains on classical hardware and runs on current quantum devices (Fig. 1). Classical simulations up to $28$ qubits validate the approach and assess its noise resilience. These show that our models can reproduce the key defining properties of each graph family, with errors that depend on feature complexity and locality.
We study the scaling by deploying $28$-, $45$-, $91$-, and $153$-qubit models on IBM’s Aachen QPU. For global features like bipartiteness, the models outperform the null-hypothesis baseline up to $45$ qubits, demonstrating their ability to reproduce the bipartite structure on real hardware; performance then declines at larger scales owing to increased noise and complexity. A relaxed measure of bipartiteness, spectral bipartivity, continues to outperform the baseline up to $91$ qubits. Notably, even at $153$ qubits, local features such as the degree distribution remain accurately captured. Finally, maximum mean discrepancy tests confirm that the learned distributions remain faithful to the targets.
@@
The paper is organized as follows. Section II defines the target graph distributions, their representations, and our features of interest. Section III describes the proposed shallow IQP model and its optimization. Section IV outlines the experimental setup, including the datasets, training, hyperparameter optimization, and sampling. Section V reports the results. Finally, section VI concludes and discusses future directions.
@@ II.2 Features of interest
These definitions measure the correlations required to reproduce any given property in an idealized setting. We now describe the features used in this work, their theoretical bodyness, and practical considerations.
@@
Density provides a coarse summary of the overall connectivity, showing whether a graph is sparse ($\rho\approx 0$) or dense ($\rho\approx 1$), but it does not reveal how edges are arranged among nodes. Since the expected density depends only on single-edge probabilities, it is a 11-bodied feature. Consequently, it should be easy to learn and reproduce.
@@
where $\mathcal{E}_{j}$ is the set of edges incident on $v_{j}$. The degree distribution of a graph is then the histogram of $\deg(v_{j})$ over all nodes $j\in\{1,\dots,M\}$. For random graphs with independent edges, the degree of a node is the sum of independent Bernoulli variables, yielding a binomial distribution. Therefore, the degree distribution is also 11-bodied. Although it has the same bodyness as the density, it should be harder to reproduce in practice, as it encodes additional information beyond the mean edge probability.
The last feature we consider is the bipartite property, which indicates whether a graph is 22-colorable—i.e., has no odd-length cycles. The degree of correlations required to determine this property increases with graph size; for an $M$-node graph, one must capture all odd-cycle correlations up to length MM—or M1M-1 if $M$ is even. Consequently, bipartiteness is a global feature, as it depends on correlations that span the entire graph.
@@ III IQP circuit born machines
$\mathbb{E}_{\mathbf{z}\sim q_{\bm{\theta}}}[f(\mathbf{z})]-\mathbb{E}_{\mathbf{z}\sim p}[f(\mathbf{z})]<\varepsilon,$ (5)
@@ III.1 Shallow IQP circuits
The foundation of our models is a NN-qubit shallow IQP circuit, building on the general definition from Nakata et al. [nakataDiagonalQuantumCircuits2014].
@@
    Diagonal Evolution: A constant depth parameterized block of gates $D_{Z}(\bm{\theta})$ that is diagonal in the Pauli-Z basis. This block is composed of single- and two-qubit rotations generated by Pauli-Z operators.
@@
The classical simulability of IQP circuits depends on the type and connectivity of the gates employed. Circuits composed solely of single-qubit rotations are trivially simulable, as they generate no entanglement. Furthermore, circuits restricted to nearest-neighbor two-qubit gates can also be efficiently simulated [fujiiQuantumCommutingCircuits2017]. Notably, circuits that combine these two layouts fall out of the simulable regime while maintaining a shallow depth [fujiiQuantumCommutingCircuits2017]. This is the chosen design for this work, one of the simplest IQP instances conjectured to be classically intractable.
This shallow ansatz maps efficiently to the target quantum processor (QPU), IBM’s Aachen, since both parameterized gates belong to the device’s native gate set. In particular, RZ gates are implemented as virtual rotations with negligible latency and error [mckayEfficientZGatesQuantum2017]. Although Hadamard gates are not native, the compiler decomposes them into RZ($\pi/2$) and SX gates, introducing only a minimal depth overhead.
@@ III.2 Optimization via the maximum mean discrepancy
$\begin{split}MMD^{2}(p,q_{\theta})&=\mathbb{E}_{\bm{x}\sim p,\bm{y}\sim p}[k(\bm{x},\bm{y})]\\ &\quad-2\mathbb{E}_{\bm{x}\sim p,\bm{y}\sim q_{\theta}}[k(\bm{x},\bm{y})]\\ &\quad+\mathbb{E}_{\bm{x}\sim q_{\theta},\bm{y}\sim q_{\theta}}[k(\bm{x},\bm{y})],\end{split}$ (8)
where $k(\bm{x},\bm{y})$ is a positive definite kernel. If $k$ is characteristic, then $MMD^{2}=0$ if and only if $p=q_{\theta}$ [JMLR:v13:gretton12a]. The three terms respectively measure intra-data similarity, model-data cross-similarity, and intra-model similarity.
@@
$MMD^{2}(\bm{\theta})=\mathbb{E}_{\bm{a}\sim\mathcal{P}{\sigma}(\bm{a})}\left[(\langle Z_{\bm{a}}\rangle_{p}-\langle Z_{\bm{a}}\rangle_{q_{\bm{\theta}}})^{2}\right],$ (10)
@@
Each bit appears independently with probability $p_{\sigma}$, so the active bit count $|\bm{a}|$ follows a binomial distribution with mean $np_{\sigma}$ [rudolphTrainabilityBarriersOpportunities2024].
@@
IV Experimental setup
This section describes the experimental framework used to evaluate the proposed workflow (Fig. 1). We begin by outlining the generation of synthetic graph datasets employed for training and validation. The following subsections detail the model training procedure, hyperparameter optimization (HPO), and sampling strategy. We then introduce the quantitative metrics used to compare the generated and target graph distributions. Together, these components form a reproducible pipeline for assessing the model’s performance.
All the code and data supporting this work are available in the following GitHub repository.
@@ IV.1 Datasets
We generated synthetic datasets for the two graph families introduced in Sec. II: Erdős–Rényi (ER) graphs, where each edge is included independently with probability $\rho$, and bipartite (BP) graphs, where a 22-colorability constraint is imposed. We use the NetworkX library [SciPyProceedings_11] to build them, ensuring that each sample represents a unique, non-isomorphic graph using the procedure described in Sec. II.1.
@@ IV.3 Hyperparameter optimization
To ensure a consistent basis for comparison across models, we perform HPO using the Optuna library [optuna_2019]. Each configuration is evaluated via repeated $k$-fold cross-validation, with two repetitions and between three and five folds depending on the dataset size.
We optimize three key parameters: the learning rate, which controls the optimizer step size; the bandwidth multiplier, which scales the median-heuristic kernel width; and the initialization multiplier, which adjusts the amplitude of the initial weights. Other settings—such as the number of ansatz layers and the optimizer choice—showed negligible influence on performance in preliminary tests and were therefore kept fixed.
@@ IV.4 Sampling
Importantly, we don’t employ any error mitigation technique or classical post-processing, which allows us to directly assess the raw performance of the quantum hardware.

### v3

# Shallow instantaneous quantum polynomial-time circuits for generative modeling on noisy intermediate-scale quantum hardware

@@ Abstract
Generative modeling is one of the most promising applications of quantum machine learning, yet training and deploying Quantum Generative Models (QGMs) on near-term hardware remains effectively intractable due to prohibitive gradient estimation and implementation costs. We propose a resource-efficient approach based on shallow Instantaneous Quantum Polynomial-time (IQP) circuits that circumvents these bottlenecks by leveraging efficient classical training while retaining the guarantee of sampling hardness. To validate this approach, we formalize graph generation as a hierarchy of physical correlations, allowing us to map abstract data features—such as edge density and bipartiteness—directly to the quantum observables required to learn them. We validate our protocol through demonstrations both on real hardware (from $28$ to $153$ qubits) and simulations ($28$ qubits). Results show that while global structural features exhibit significant degradation beyond $91$ qubits, our models achieve high-precision reproduction of local correlations, even up to $153$ qubits. These findings establish shallow IQP circuits as a robust, scalable candidate for generative tasks on current Noisy Intermediate-Scale Quantum (NISQ) devices.
@@ I Introduction
Generative modeling stands as a cornerstone of unsupervised machine learning, enabling the synthesis of data that captures the underlying statistics of complex distributions. Within this domain, Quantum Generative Models (QGMs) have emerged as a promising frontier, leveraging quantum computation to represent probability distributions that are potentially intractable for classical systems [zoufalGenerativeQuantumMachine2021, swekeQuantumClassicalLearnability2021, zengLearningInferenceGenerative2019]. A particularly compelling architecture in this landscape is the Quantum Circuit Born Machine (QCBM) [coyleBornSupremacyQuantum2020, mitaraiQuantumCircuitLearning2018a, liuDifferentiableLearningQuantum2018a, kissConditionalBornMachine2022]. However, training QCBMs on near-term hardware presents significant hurdles, primarily due to the poor scaling of gradient estimation costs and the prevalence of barren plateaus, which currently restricts training to small-scale models. Furthermore, even in instances where large-scale models can be trained without barren plateaus, efficient implementation and robust guarantees of classical hardness remain elusive [cerezoDoesProvableAbsence2025].
In this work, we propose a novel generative modeling approach that overcomes these bottlenecks by leveraging the unique properties of Instantaneous Quantum Polynomial-time (IQP) circuits. IQP circuits occupy a prime position for near-term applications: they are conjectured to be hard to sample from classically [shepherdInstantaneousQuantumComputation2009, marshallImprovedSeparationQuantum2024], yet their specific algebraic structure permits the efficient classical calculation of certain output statistics that can be used for training [recio-armengolTrainClassicalDeploy2025, kastureProtocolsClassicallyTraining2023]. We introduce a hybrid workflow based on shallow IQP circuits that exploits this duality—training the model on classical hardware and deploying it on quantum processors for sampling [recio-armengolTrainClassicalDeploy2025, recio-armengolIQPoptFastOptimization2025]—while minimizing the resources required for implementation and retaining classical hardness. This approach effectively circumvents the scalability and implementation challenges inherent in standard QCBMs.
@@
To rigorously validate this approach and characterize the expressivity of shallow IQP circuits, we use graph generation as a testbed. Graphs represent relationships between entities and provide a unified framework to describe complex systems via their components and interactions [newmanNetworks2018]. Their versatility has driven the development of graph generation methods for applications ranging from drug discovery to scheduling [drobyshevskiyRandomGraphModeling2020, bonifatiGraphGeneratorsState2021, bonginiMolecularGenerativeGraph2021, yangMoleculeGenerationDrug2024, cordeiroRandomGraphGeneration2010]. Learning a graph distribution remains challenging because node interactions span both local and global dependencies [groverGraphiteIterativeGenerative2019, tranDeepNCDeepGenerative2020, bacciuEdgebasedSequentialGraph2020, simonovskyGraphVAEGenerationSmall2018, flam-shepherdGraphDeconvolutionalGeneration2020, niuPermutationInvariantGraph2020]. As the number of nodes increases, the combinatorial complexity of pairwise and higher-order relationships grows rapidly: adding a single node to an $M$-node graph introduces $M$ new potential edges, so that the number of possible configurations increases from $2^{\binom{M}{2}}$ to $2^{\binom{M}{2}}\cdot 2^{M}$. This exponential scaling implies that even modest increases in graph size can dramatically expand the hypothesis space. Although practical problems typically restrict attention to structured graph families, models must still possess sufficient expressivity to navigate this high-dimensional space [guoSystematicSurveyDeep2022, simonovskyGraphVAEGenerationSmall2018].
Rather than treating graph generation as a downstream application, we utilize it to map abstract data features to physical correlations. Simple properties like edge density correspond to local ($1$-body) correlations, while structural constraints like bipartiteness correspond to global ones. This formalization allows us to disentangle the model’s capacity to learn features of different complexities.
We demonstrate the efficacy of this approach by training shallow IQP models to generate random bipartite and Erdős–Rényi graph distributions. Our demonstration results, obtained on an IBM Heron r2 $156$ qubit processor, show that the proposed models are highly effective at reproducing low-bodied correlations, even up to $153$ qubits. While global structural features exhibit higher performance degradation beyond $91$ qubits, due to hardware noise. These findings establish shallow IQP circuits as a resource-efficient candidate for near-term generative tasks.
@@
The paper is organized as follows. Section II defines the target graph distributions, their representations, and our features of interest. Section III describes the proposed shallow IQP model and its optimization. Section IV outlines the demonstration setup, including the datasets, training, hyperparameter optimization, and sampling. Section V reports the results. Finally, section VI concludes and discusses future directions.
@@ II.2 Features of interest
These definitions measure the correlations required to learn to reproduce any given property in an idealized setting. We now describe the features used in this work, their theoretical bodyness, and practical considerations.
@@
Density provides a coarse summary of the overall connectivity, showing whether a graph is sparse ($\rho\approx 0$) or dense ($\rho\approx 1$), but it does not reveal how edges are arranged among nodes. Since the expected density depends only on single-edge probabilities, it is a $1$-bodied feature. Consequently, it should be easy to learn and reproduce.
@@
where $\mathcal{E}_{j}$ is the set of edges incident on $v_{j}$. The degree distribution of a graph is then the histogram of $\deg(v_{j})$ over all nodes $j\in\{1,\dots,M\}$. For random graphs with independent edges, the degree of a node is the sum of independent Bernoulli variables, yielding a binomial distribution. Therefore, the degree distribution is also $1$-bodied. Although it has the same bodyness as the density, it should be harder to reproduce in practice, as it encodes additional information beyond the mean edge probability.
Another feature we consider is the bipartite property, which indicates whether a graph is $2$-colorable—i.e., has no odd-length cycles. The degree of correlations required to determine this property increases with graph size; for an $M$-node graph, one must capture all odd-cycle correlations up to length $M$, or M1M-1 if $M$ is even. Consequently, bipartiteness is a global feature, as it depends on correlations that span the entire graph.
@@ III IQP circuit born machines
$\left|\mathbb{E}_{\mathbf{z}\sim q_{\bm{\theta}}}[f(\mathbf{z})]-\mathbb{E}_{\mathbf{z}\sim p}[f(\mathbf{z})]\right|<\varepsilon,$ (5)
@@ III.1 Shallow IQP circuits
The foundation of our models is a $N$-qubit shallow IQP circuit, building on the general definition from Nakata et al. [nakataDiagonalQuantumCircuits2014].
@@
    Diagonal Evolution: A parameterized unitary block $D_{Z}(\bm{\theta})$ diagonal in the computational basis. This block consists of a sequence of commuting rotation gates generated by Pauli-$Z$ strings acting on subsets of qubits $S$:
    $D_{Z}(\bm{\theta})=\prod_{S\in\mathcal{I}}\exp\left(-i\theta_{S}\prod_{j\in S}Z_{j}\right),$ (6)
    where $\mathcal{I}$ is a collection of subsets of $\{1,\dots,N\}$ specifying the interactions. The collection $\mathcal{I}$ is chosen such that the total circuit depth is $\mathcal{O}(1)$, independent of the number of qubits $N$.
@@
The classical simulability of IQP circuits depends strictly on the connectivity and type of gates employed, rather than circuit depth, as the commuting nature of the gates renders the temporal order of operations irrelevant. Circuits composed solely of single-qubit rotations are trivially simulable, as they generate no entanglement. Furthermore, circuits restricted to nearest-neighbor two-qubit gates on a planar grid can also be efficiently simulated, regardless of the number of layers. This is because their output probability distribution maps to the partition function of a 2D Ising model with zero magnetic field, a problem exactly solvable via the Pfaffian method (free fermions) [fujiiQuantumCommutingCircuits2017].
Notably, combining these two simulable designs in a planar 2D layout pushes the model out of the simulable regime [fujiiQuantumCommutingCircuits2017]. Introducing single-qubit rotations corresponds to adding magnetic fields to the underlying Ising model. This modification breaks the free-fermion mapping and renders the calculation of the partition function #P-hard [fujiiQuantumCommutingCircuits2017]. This combination—nearest-neighbor interactions plus single-qubit rotations—constitutes the design chosen for this work, ensuring the ansatz remains shallow yet conjectured to be classically intractable.
This ansatz maps efficiently to the target quantum processor (QPU), IBM’s Aachen, since both parameterized gates belong to the device’s native gate set. In particular, RZ gates are implemented as virtual rotations with negligible latency and error [mckayEfficientZGatesQuantum2017]. Although Hadamard gates are not native, the compiler decomposes them into RZ($\pi/2$) and SX gates, introducing only a minimal depth overhead.
@@ III.2 Optimization via the maximum mean discrepancy
$\begin{split}\text{MMD}^{2}(p,q_{\theta})&=\mathbb{E}_{\bm{x}\sim p,\bm{y}\sim p}[k(\bm{x},\bm{y})]\\ &\quad-2\mathbb{E}_{\bm{x}\sim p,\bm{y}\sim q_{\theta}}[k(\bm{x},\bm{y})]\\ &\quad+\mathbb{E}_{\bm{x}\sim q_{\theta},\bm{y}\sim q_{\theta}}[k(\bm{x},\bm{y})],\end{split}$ (9)
where $k(\bm{x},\bm{y})$ is a positive definite kernel. If $k$ is characteristic, then $\text{MMD}^{2}=0$ if and only if $p=q_{\theta}$ [JMLR:v13:gretton12a]. The three terms respectively measure intra-data similarity, model-data cross-similarity, and intra-model similarity.
@@
$\text{MMD}^{2}(\bm{\theta})=\mathbb{E}_{\bm{a}\sim\mathcal{P}{\sigma}(\bm{a})}\left[(\langle Z_{\bm{a}}\rangle_{p}-\langle Z_{\bm{a}}\rangle_{q_{\bm{\theta}}})^{2}\right],$ (11)
@@
Each bit appears independently with probability $p_{\sigma}$, so the active bit count $|\bm{a}|$ follows a binomial distribution with mean $np_{\sigma}$ [rudolphTrainabilityBarriersOpportunities2024]. A derivation of the MMD as a quantum observable is shown in the Appendix A.
@@
IV Demonstration setup
This section describes the demonstration framework used to evaluate the proposed workflow (Fig. 1). We begin by outlining the generation of synthetic graph datasets employed for training and validation. The following subsections detail the model training procedure, hyperparameter optimization (HPO), and sampling strategy. We then introduce the quantitative metrics used to compare the generated and target graph distributions. Together, these components form a reproducible pipeline for assessing the model’s performance.
All the code and data supporting this work are openly available in Zenodo [ballo_gimbernat_2026_zenodo].
@@ IV.1 Datasets
We generated synthetic datasets for the two graph families introduced in Sec. II: Erdős–Rényi (ER) graphs, where each edge is included independently with probability $\rho$, and bipartite (BP) graphs, where a $2$-colorability constraint is imposed. We use the NetworkX library [SciPyProceedings_11] to build them, ensuring that each sample represents a unique, non-isomorphic graph using the procedure described in Sec. II.1.
@@ IV.3 Hyperparameter optimization
To ensure a consistent basis for comparison across models, we perform HPO using the Optuna library [optuna_2019] employing a Tree-structured Parzen Estimator (TPE) sampler [watanabeTreeStructuredParzenEstimator2025]. We allocate a budget of $60$ trials per dataset. This allocation is justifiable given that our search space has only three hyperparameters. Even under the conservative theoretical bounds of random search, $60$ iterations provide a $95\%$ probability ($1-0.95^{60}\approx 0.954$) of sampling a configuration within the top $5\%$ of the search space volume [bergstraRandomSearchHyperParameter]. Since TPE leverages Bayesian optimization to prioritize promising regions, it is expected to locate optimal configurations more efficiently than this baseline. Furthermore, each configuration is evaluated via repeated $k$-fold cross-validation, with three repetitions and five folds. Once the best configuration is found, it is trained three separate times, with small perturbations on the initialization, and the best performing model is saved for sampling.
We optimize three key parameters: the learning rate, which controls the optimizer step size; the bandwidth multiplier, which scales the median-heuristic kernel width; and the initialization multiplier, which adjusts the amplitude of the initial weights. Other settings—such as the optimizer choice—showed negligible influence on performance in preliminary tests and were therefore kept fixed.
@@ IV.4 Sampling
Importantly, we do not employ any error mitigation technique or classical post-processing, which allows us to directly assess the raw performance of the quantum hardware.